# Reasoning or Reciting? Exploring the Capabilities and Limitations of Language Models Through Counterfactual Tasks

### v1

@@ 5.6 Qualitative Analysis of Drawing Results
Figure 7: Visualizations of objects drawn by GPT-4 under the default (upright) and counterfactual conditions: vertical flip (Vflip), rotates 90 degrees (R90), and 180 degrees (R180). In all cases, the CCC (not shown) passes. We show the original output, without flipping/rotating back as in our quantitative evaluation (§A.6). For the counterfactual settings, GPT-4 either does not transform the objects as instructed (e.g., house and penguin) or struggles to draw meaningful objects (e.g., cake and unicorn).
@@ 6 Discussion
Why do we care about counterfactual worlds? Wouldn’t a model for only the default task instantiation be nonetheless useful? It is certainly true that such a model would still be useful. However, many of the counterfactual worlds that we investigate actually are not very distant so that model performance under them still bears utility. For example, addition in different bases is certainly useful for many applications. More generally, we are necessarily interested in the counterfactual tasks themselves; we are only interested in them insofar as performance on these tasks can serve as a measurable proxy for the generalizability of these models and their underlying reasoning capabilities.
@@
Can some more carefully designed prompts eliminate the default-counterfactual gap? This is always a possibility, and one that we can never tractably rule out. Nevertheless, given the consistent patterns across our tasks, we believe that a prompt that completely bridges the default-counterfactual gap is unlikely. Our in-context learning experiment (§5.5) shows that this gap could be reduced, but not fully removed, by more informative prompts. It would be interesting to apply more advanced prompting techniques (Wang et al., 2023a, 2022a; Yao et al., 2023; Sordoni et al., 2023; i.a.) to our counterfactual tasks. We considered 0-shot chain-of-thought in this work, which did not fully bridge the default-counterfactual gap, but we leave the exploration of these more recent prompting techniques to future work.
@@ 7.1 Underestimation
Retrieving notes in melodies in different keys faces a similar issue: one way of retrieving a note in a melody in an uncommon key would be to first retrieve it in a canonical key and then transpose it to the desired key. With this strategy, the counterfactual task consists of 2 steps and is harder than (and requires first) completing the 1-step original task. This strategy is not the only way to solve the task: an alternate one would be to recall a melody as a series of abstract relations in a scale and directly map them onto notes in a target key. However, the 2-step process is a natural one that is often employed by musicians. The counterfactual setup thus introduces a confounder: low performance may be driven by the increased difficulty of the counterfactual task, rather than overfitting to melodies in their canonical keys, if models are employing two-step strategy. However, since both strategies are available to models and we do not prompt them to use a particular one, reliance on this two-step strategy may itself be indicative of overfitting to the original canonical keys.
@@ 8 Related Work
Much prior work has investigated the extent to which LMs acquire a grounded understanding of the world through text-only training (Piantadosi and Hill, 2022; Zhang et al., 2020; Ilharco et al., 2021; Li et al., 2021; i.a.). These works have generally found that conceptual structures of certain concepts (e.g., color, size) often plausibly mirror those of the grounded world (Abdou et al., 2021; Patel and Pavlick, 2022; Mollo and Millière, 2023). As in our study, these studies are a test of generalization—such structures would not manifest if the concepts were memorized in a one-hot-like manner. But our evaluation differs in that it targets the reasoning process instead of the generalization to new concepts or conceptual structures (Kondo et al., 2023). While prior work identified that the latter is embedded in LMs, we found that they do not fully learn the former.
@@
Our counterfactual perturbations can be informally viewed as interventions under a causal inference framework (Pearl, 2009). This relationship has been explored in machine learning and NLP for commonsense reasoning (Kıcıman et al., 2023), interpretability (Elazar et al., 2021; Geiger et al., 2021, 2022), spurious correlation detection (Veitch et al., 2021; Eisenstein, 2022), and fairness (Kusner et al., 2017; Nabi and Shpitser, 2018). Under this perspective, the failure of generalization to counterfactual worlds that we observe in LMs can be viewed as a failure to robustly learn the causal effects of world states on our evaluated tasks.
@@ 9 Conclusion
Through our counterfactual evaluation on 11 tasks, we identified a consistent and substantial degradation of LM performance under counterfactual conditions. We attribute this gap to overfitting to the default task variants, and thus encourage future LM analyses to explicitly consider abstract task ability as detached from observed task performance, especially when these evaluated task variants might exist in abundance in the LM pretraining corpora.
@@ A.3 Basic Syntactic Reasoning
We designed the CCC to assess how well LMs understand the instruction that explains the difference of word orders in the counterfactual settings. We synthetically generate 100 simple three-word sentences (e.g., anna saw john) in five counterfactual word orders (e.g., anna john saw in SOV), and ask LMs to reconstruct the original English sentences in SVO order. Conceptually, this is equivalent to asking the model to identify the subject, verb, and object in the perturbed order, but using a format that is more familiar to the LM.
@@ A.6 Drawing
For the counterfactual settings, we ask the LMs to draw the same object, but vertically flipped, or rotated by 90°or 180°. We also ask the LMs to avoid using any transformation functions such as rotate and scale to avoid shortcuts. Before our quantitative evaluation, we flip/rotate back the generated drawing.
@@ Appendix C Raw Results
PaLM-2 016.0 05.0 02.0 1.0 00.0 0.0
Table 34: Breakdown for the SET game test results (w/ 0-CoT) when the model needs to find different number of cards (c) in a SET as analyzed in §5 (in accuracy; %).

### v3

@@ 5.6 Qualitative Analysis of Drawing Results
Figure 7: Visualizations of objects drawn by GPT-4 under the default (upright) and counterfactual conditions: vertical flip (Vflip, i.e. upside-down), rotates 90 degrees (R90), and 180 degrees (R180). In all cases, the CCC (not shown) passes. We show the original output, without flipping/rotating back as in our quantitative evaluation (§A.6). For the counterfactual settings, GPT-4 either does not transform the objects as instructed (e.g., house and penguin) or struggles to draw meaningful objects (e.g., cake and unicorn).
@@ 6 Discussion
Why do we care about counterfactual worlds? Wouldn’t a model for only the default task instantiation be nonetheless useful? It is certainly true that such a model would still be useful. However, many of the counterfactual worlds that we investigate are not very distant so that model performance under them still bears utility. For example, addition in different bases is certainly useful for many applications. More generally, we are necessarily interested in the counterfactual tasks themselves; we are only interested in them insofar as performance on these tasks can serve as a measurable proxy for the generalizability of these models and their underlying reasoning capabilities.
@@
Can some more carefully designed prompts eliminate the default-counterfactual gap? This is always a possibility, and one that we can never tractably rule out. Nevertheless, given the consistency of the gap across our tasks (which use different prompts) and the 0-shot CoT setting, we believe that a prompt that completely bridges the default-counterfactual gap is unlikely. Our in-context learning experiment (§5.5) further shows that while this gap could be reduced by more informative prompts, it is not fully removed. It would be interesting to apply more advanced prompting techniques (Wang et al., 2023a, 2022a; Yao et al., 2023; Sordoni et al., 2023; i.a.) to our counterfactual tasks. We considered 0-shot chain-of-thought in this work, which did not fully bridge the default-counterfactual gap, but we leave the exploration of these more recent prompting techniques to future work.
@@ 7.1 Underestimation
Retrieving notes in melodies in different keys faces a similar issue. We expect similar retrieval difficulty under different keys if the model recalls a melody as a series of abstract relations in a scale and directly maps them onto notes in a target key. However, an alternative strategy would be to first retrieve the note in a canonical key and then transpose it to the desired uncommon key. This 2-step process is a natural one that is often employed by musicians. And with this strategy, the counterfactual task consists of 2 steps and is harder than (and requires first) completing the 1-step original task. The counterfactual setup thus introduces a confounder: low performance may be driven by the increased difficulty of the counterfactual task, rather than overfitting to melodies in their canonical keys, if models are employing two-step strategy. However, since both strategies are available to models and we do not prompt them to use a particular one, reliance on this two-step strategy may itself be indicative of overfitting to the original canonical keys.
@@ 8 Related Work
Much prior work has investigated the extent to which LMs acquire a grounded understanding of the world through text-only training (Piantadosi and Hill, 2022; Zhang et al., 2020; Ilharco et al., 2021; Li et al., 2021; i.a.). These studies have generally found that conceptual structures of certain concepts (e.g., color, size) often plausibly mirror those of the grounded world (Abdou et al., 2021; Patel and Pavlick, 2022; Mollo and Millière, 2023). As in our study, these studies are a test of generalization—such structures would not manifest if the concepts were memorized in a one-hot-like manner. But our evaluation differs in that it targets the reasoning process instead of the generalization to new concepts or conceptual structures (Kondo et al., 2023). While prior work identified that the latter is embedded in LMs, we found that they do not fully learn the former.
@@
Our counterfactual perturbations can be informally viewed as interventions under a causal inference framework (Pearl, 2009). This relationship has been explored in machine learning and NLP for commonsense reasoning (Kıcıman et al., 2023), interpretability (Elazar et al., 2021; Geiger et al., 2021, 2022), spurious correlation detection (Veitch et al., 2021; Eisenstein, 2022), fairness (Kusner et al., 2017; Nabi and Shpitser, 2018), etc. Under this perspective, the failure of generalization to counterfactual worlds that we observe in LMs can be viewed as a failure to robustly learn the causal effects of world states on our evaluated tasks.
@@ 9 Conclusion
Through our counterfactual evaluation on 11 tasks, we identified consistent and substantial degradation of LM performance under counterfactual conditions. We attribute this gap to overfitting to the default task variants, and thus encourage future LM analyses to explicitly consider abstract task ability as detached from observed task performance, especially when these evaluated task variants might exist in abundance in the LM pretraining corpora. Furthermore, insofar as this degradation is a result of the LMs’ being trained only on surface form text, it would also be interesting future work to see if more grounded LMs (grounded in the “real” world, or some semantic representation, etc.) are more robust to task variations.
@@ A.3 Basic Syntactic Reasoning
We designed the CCC to assess how well LMs understand the instruction that explains the difference of word orders in the counterfactual settings. We synthetically generate 100 simple three-word sentences (e.g., “anna saw john”) in five counterfactual word orders (e.g., “anna john saw” in SOV), and ask LMs to reconstruct the original English sentences in SVO order. Conceptually, this is equivalent to asking the model to identify the subject, verb, and object in the perturbed order, but using a format that is more familiar to the LM.
@@
A note on CCC results.
The results for this task are shown in Table 22. Generally, the models pass our crafted CCC challenge with decent accuracy, but we observed that, in a few cases, the LMs are confused by the reconstruction ambiguity explained above. GPT-3.5 and Claude fail in the OVS settings where they often directly copy the original sentence—e.g., instead of reconstructing “anna saw john” to “john saw anna”, they simply copy the original sentence “anna saw john” as the output. Similarly, PaLM-2 often incorrectly reverses the subject and object in the SOV and VSO settings—e.g., instead of reconstructing “calls tom lucas” to “tom calls lucas”, it outputs “lucas calls tom”.
@@ A.6 Drawing
For the counterfactual settings, we ask the LMs to draw the same object, but vertically flipped (i.e., upside-down), or rotated by 90°or 180°. We also ask the LMs to avoid using any transformation functions such as rotate and scale to avoid shortcuts. Before our quantitative evaluation, we flip/rotate back the generated drawing.
@@ Appendix C Raw Results
PaLM-2 062.0 13.0 10.0 1.0 03.0 1.0
Table 34: Breakdown for the SET game test results (with 0-CoT) when the model needs to find different number of cards (c) in a SET as analyzed in §5 (in accuracy; %).